AnimeGhoul8863:

LAST JUAN! Carl can paint a room 3 hours faster than Jennifer can. If they work together, they can complete the job in 2 hours. Using complete sentences, explain each step in figuring out how to determine the time it would take Jennifer to complete this job on her own.

AnimeGhoul8863:

@Hero e.e

Hero:

Use the formula

Hero:

I know you have the formula. I gave it to you. So at least write that out first.

AnimeGhoul8863:

ok hold on

AnimeGhoul8863:

\[t=\frac{ CJ }{ C+J }\]

Hero:

Okay wait.

AnimeGhoul8863:

e.e

Hero:

This is a bit more involved than you think. The only given value is \(t\)

Hero:

It says Carl can paint a room 3 hours faster than Jennifer can. So we don't have any given values for \(C\) or \(J\)

AnimeGhoul8863:

ok

Hero:

However we can create a statement based on this: "Carl can paint a room 3 hours faster than Jennifer".

Hero:

It translates to \(C = J + 3\)

AnimeGhoul8863:

ok

Hero:

In other words we insert \(J + 3\) in place of \(C\) in the formula

AnimeGhoul8863:

so \[\frac{ (J+3)(J) }{ ?}\]

AnimeGhoul8863:

like this?

Hero:

You're on the right track. Finish it. Make sure you set it equal to \(t\) for consistency.

AnimeGhoul8863:

e.e hold on i just got lost

AnimeGhoul8863:

how am i finishing this?

Hero:

How did you get lost. You were replacing \(C\) with the expression in the formula

AnimeGhoul8863:

yes ok so C= J+3 now what this is where i got lost

Hero:

You did the numerator correctly. All you have to do it do the denominator.

AnimeGhoul8863:

\[\frac{ (J+3)(J) }{ (J+3)+(J)}\]

AnimeGhoul8863:

like this??

Hero:

Exactly but set that equal to \(t\)

AnimeGhoul8863:

\[t=\frac{ (J+3)(J) }{ (J+3)+(J)}\]

Hero:

Precisely

AnimeGhoul8863:

so now what

Hero:

Now insert the given value for \(t\)

Hero:

Do you know what it is? Do you remember what \(t\) represents? Go back to the previous question if you don't remember.

AnimeGhoul8863:

t represents time it takes to finish there objective

Hero:

Is that amount given in the problem?

AnimeGhoul8863:

ummmmmmmm idk let me check the problem

AnimeGhoul8863:

tbh idk and i dont think it is

Hero:

Go to the previous problem and find out what \(t\) represents

AnimeGhoul8863:

i just said t represents the time it takes to finish there objective as in painting

Hero:

And what is the equivalent of that here? What's another word for "finish" in the problem?

AnimeGhoul8863:

complete e.e

Hero:

Bingo. So do you have it now? What is the value of \(t\)?

AnimeGhoul8863:

ummmmmm not really so the value of t is c?

Hero:

Read the problem to yourself once more. I'm not going to give this to you since it is so obvious.

AnimeGhoul8863:

oh wait so t=2

Hero:

Correct

AnimeGhoul8863:

ok but what does that have to do with the time it takes Jennifer to do it her own

Hero:

We will solve for \(J\). If you notice, the equation is all in terms of \(J\) now.

AnimeGhoul8863:

\[2=\frac{ (J+3)(J) }{ (J+3)+(J) }\]

AnimeGhoul8863:

So it looks like this????

Hero:

Yes, but the next step is to add the J's in the denominator.

AnimeGhoul8863:

ok......

AnimeGhoul8863:

would it be 2J+3????

Hero:

Correct

AnimeGhoul8863:

so it would look like \[2=\frac{ (J+3)(J) }{ 2J+3 }\]

AnimeGhoul8863:

@Hero what next

Hero:

Multiply both sides by the denominator

AnimeGhoul8863:

ok so after doing that would it equal 2J+3 = J(J+3)

Hero:

Distribute \(J(J + 3)\)

AnimeGhoul8863:

it would equal \[J^2+3J\]

AnimeGhoul8863:

i think atleast

Hero:

Yes that is correct but remember, for consistency include what is on the other side of the equal sign.

Hero:

Because the next step is to subtract \(2J\) from both sides

AnimeGhoul8863:

wait so \[2=\frac{ (J+3)(J) }{ J^2+3J }\] this is what it is

AnimeGhoul8863:

now we subtract 2j from both sides?

Hero:

You're going backwards. You already multiplied both sides by the denominator? Remember?

Hero:

Do you not remember being at this step: \(2J+3 = J(J+3)\) We're not dealing with a fraction anymore

Hero:

The next step was to distribute the expression on the RHS

AnimeGhoul8863:

ohhhhhhhhhh ok thats where i got lost i thought we where still doing that

Hero:

Do you know where we are now?

Hero:

If so, please post it

AnimeGhoul8863:

all i can think of is it would equal 4J+6 but i think i could be wrong

Hero:

We were here: \(2J+3 = J(J+3)\) The next step was to distribute \(J(J + 3)\)

Hero:

It's not a good idea to watch Hanna Montana at the same time you're doing math btw

AnimeGhoul8863:

e.e what

Hero:

Nevermind

Hero:

I'm just saying it is best not to be distracted while doing this

AnimeGhoul8863:

anyway i already did distribute it e.e

AnimeGhoul8863:

i got \[J^2+3J\]

Hero:

Okay if you did, you should post that step here. Post the ENTIRE equation not just one side of it.

Hero:

Look at my previous step if you need help.

AnimeGhoul8863:

wait what your previous step wait hold on either your going to fast or im getting lost again let me read for a second

Hero:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Hero We were here: \(2J+3 = J(J+3)\) The next step was to distribute \(J(J + 3)\) \(\color{#0cbb34}{\text{End of Quote}}\)

AnimeGhoul8863:

I ALREADY DISTRIBUTED IT LIKE 3 TIMES e.e

Hero:

Yes, but show both sides of the equation

AnimeGhoul8863:

Both sides of the equation?

Hero:

Showing both sides is confirmation you know what's going on

Hero:

YES

AnimeGhoul8863:

ok hold on

AnimeGhoul8863:

i officially got nothing i have no idea what to do

Hero:

We began with this \[t=\frac{ CJ }{ C+J }\] Then you figured out the expression for \(C\): \[t=\frac{ (J+3)(J) }{ (J+3)+(J)}\] Next you figured out the value of \(t\): \[2=\frac{ (J+3)(J) }{ 2J+3 }\] Then you multiplied both sides by the denominator to get: \(2(2J + 3) = J(J + 3)\)

Hero:

Notice the whole time we're dealing with expressions on both sides of the equal sign.

AnimeGhoul8863:

yes

Hero:

Do you want me to also hightlight them in colors? At each step, you perform a step while showing both sides of the equal time.

Hero:

equal sign*

Hero:

I don't know how you got confused when we were going step by step. You seemed to be following.

AnimeGhoul8863:

i am following the best i can but its not easy for me to understand this

AnimeGhoul8863:

We began with this \[t=\frac{ CJ }{ C+J }\] Then you figured out the expression for \(C\): \[t=\frac{ (J+3)(J) }{ (J+3)+(J)}\] Next you figured out the value of \(t\): \[2=\frac{ (J+3)(J) }{ 2J+3 }\] Then you multiplied both sides by the denominator to get: \(2(2J + 3) = J(J + 3)\) Then we distributed and it equaled \[J^2+ 3J\]

AnimeGhoul8863:

this is where i am

Hero:

Right, so all you had to do is post the other side of the equation: \(2(2J + 3) = J^2 + 3J\) I don't think it was difficult for you to realize what I was asking you to do. Post both sides of the equal sign at every step.

AnimeGhoul8863:

thats not what i thought you ment at all tbh sorry i thought you ment something different

Hero:

Now you have to distribute the expression other side of the equation: \(2(2J + 3)\)

AnimeGhoul8863:

We began with this \[t=\frac{ CJ }{ C+J }\] Then you figured out the expression for \(C\): \[t=\frac{ (J+3)(J) }{ (J+3)+(J)}\] Next you figured out the value of \(t\): \[2=\frac{ (J+3)(J) }{ 2J+3 }\] Then you multiplied both sides by the denominator to get: \(2(2J + 3) = J(J + 3)\) Then we distributed and it equaled \[J^2+3J\] Now we have \[2(2J+3)=J^2+3J\] we must distribute again but this time on the other side \[2(2J+3)=4J+6\]

AnimeGhoul8863:

like this?

Hero:

Except you've now eliminated the other side of the equation that contains the \(J^2\)

Hero:

\(4J + 6\) is correct though but that is only one side of the equation

Hero:

You must include the other side of the original equation in the next step

Hero:

It should be \(4J + 6 = J^2 + 3J\)

Hero:

I've never seen anyone get confused on this particular thing before.

AnimeGhoul8863:

Im sorry Hero i know im hard to work with ive been told that my whole life im horrible with math it just never soaks in

AnimeGhoul8863:

anyway now lets move on

AnimeGhoul8863:

so we now have the equation as \(4J + 6 = J^2 + 3J\)

Hero:

Yes, the next step is to subtract \(4J\) from both sides

AnimeGhoul8863:

ok and that would equal \[J^2-J-6=0\] then we must solve using the quadratic formula which would give us the finishing numbers of J=3, J=-2?

Hero:

Looks good to me

Hero:

But of course you dis-regard the negative number

Hero:

As there is no such thing as negative hours

AnimeGhoul8863:

so would the answer be Jennifer takes 3/2 hours to pain the room by her self

Hero:

smh

Hero:

There's only one answer

AnimeGhoul8863:

ik i just realized what i did hold on

Hero:

Or rather, there is only one value of J

Hero:

You have one positive value and one negative value. The negative value is dis-regarded because it makes no sense in terms of what the value of J represents.

AnimeGhoul8863:

so do we put it as 3,2 or 3/2 or 2/3 as our answer

Hero:

@AnimeGhoul8863 there are two values of J. One of them we discard because the value is negative

Hero:

I don't know how many times I have to say it before you "realize"

AnimeGhoul8863:

T~T i dont purposely not understand it geez sorry so the answer is 3 got it thanks for your help

Hero:

Yes J = 3

AnimeGhoul8863:

thank you for the help i think i understand it more than i did to begin with

Hero:

Glad to hear it.